# Exploring Kolmogorov-Arnold networks for realistic image sharpness assessment

### v2

@@ IV-C Current achievement on the databases
The results of TaylorKANs using mid-level features (a) and high-level features (b) and several state-of-the-art works are shown in Table IV. These works [15, 14, 16, 17, 18, 19, 20] design novel CNNs and take images as input rather than features.
@@
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
TaylorKAN a 0.757 0.750 0.550 0.859 0.843 0.666 0.608 0.488 0.342 0.719 0.640 0.458
TaylorKAN b 0.792 0.788 0.587 0.750 0.756 0.565 0.667 0.563 0.405 0.831 0.801 0.610
[14] 0.861 0.846 0.934 0.930 0.873 0.855 0.923 0.901
[15] 0.873 0.858 0.890 0.854 0.901 0.896
[16] 0.886 0.874 0.880 0.865 0.916 0.904
[17] 0.891 0.875 0.898 0.882 0.921 0.912
[18] 0.9284 0.9128 0.9175 0.8992 0.9362 0.9209
[19] 0.9109 0.8875 0.9141 0.8947 0.9413 0.9271
[20] 0.900 0.875 0.910 0.904 0.908 0.919
A significant performance gap exists between CNNs and TaylorKAN in score prediction. CNNs achieve PLCC $\geq$ 0.86 and SRCC $\geq$ 0.84, which are notably higher than those of the proposed KAN. These works benefit not only from advanced architectures of image convolution and feature pooling but also from sophisticated module designs, such as hinge constraints on learning uncertainty [15], multi-scale spatial pyramid pooling [17], self-supervised learning [19], and multitask learning [20]. Therefore, it becomes urgent to integrate prior knowledge and advanced modules into KANs for improving representation power and score prediction performance [23, 24].
When using TaylorKAN for score prediction, deeply learned features can increase, but not substantially, metric values on three databases. On the one hand, since perceptual similarity is an emergent property shared across deep visual representation [35], it is not surprise that deep features are effective for image quality representation [21, 19]. On the other hand, selecting the most informative subsets of deep features is urgently important for KANs to enhance accuracy and efficiency in score prediction, while avoiding the challenges of high-dimensional feature processing.
@@ V Conclusions
Besides SVR and MLP, six KANs are evaluated on four realistic databases for score prediction using mid-level features and high-level features. KANs achieve superior or competitive results over SVR and MLP, showing promise on improving the performance on score prediction and related tasks.

### v3

@@ IV-C Current achievement on the databases
The results of TaylorKANs using mid-level features (a) and high-level features (b) and several state-of-the-art (SOTA) works are shown in Table IV. These works [11, 10, 12, 13, 14, 15, 16] develop novel CNNs that take images rather than features as input.
@@
BID2011 [26] CID2013 [27] CLIVE [28] KonIQ-10k [29]
PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
TaylorKAN a 0.756 0.782 0.871 0.851 0.668 0.582 0.766 0.699
TaylorKAN b 0.797 0.813 0.788 0.780 0.696 0.598 0.850 0.811
SARQUE [10] 0.861 0.846 0.934 0.930 0.873 0.855 0.923 0.901
UNIQUE [11] 0.873 0.858 0.890 0.854 0.901 0.896
REQA [12] 0.886 0.874 0.880 0.865 0.916 0.904
CSPP-IQA [13] 0.891 0.875 0.898 0.882 0.921 0.912
StairIQA [14] 0.9284 0.9128 0.9175 0.8992 0.9362 0.9209
QPT-ResNet50 [15] 0.9109 0.8875 0.9141 0.8947 0.9413 0.9271
LIQE [16] 0.900 0.875 0.910 0.904 0.908 0.919
A significant performance gap is found between the SOTA works and TaylorKAN in score prediction. The CNNs achieve PLCC $\geq$ 0.86 and SRCC $\geq$ 0.84, which are much higher than those of the evaluated models (Table II and III). These SOTA works benefit not only from advanced architectures of image convolution and feature pooling but also from sophisticated module designs, such as uncertainty learning [11], spatial pooling [13], self-supervised learning [15], and multitask learning [16]. Thus, one way to improve representation power and score prediction performance is to integrate prior knowledge and advanced modules into KANs [20]. When using TaylorKAN for score prediction, high-level features can increase, but not substantially, metric values on three databases. On the one hand, the perceptual similarity is an emergent property shared across deep visual embedding [31] and thus, it is not surprise that deep features are effective for image quality representation [15]. On the other hand, how to figure out the most informative subsets of deep features is important for KANs to enhance effectiveness and efficiency in score prediction, while avoiding the challenges of high-dimensional feature processing.
This study has several limitations. Firstly, a broader range of informative features, potentially numbering in the hundreds or thousands, could be collected and refined by selecting the most relevant ones. These selected features can then be applied in KANs for further data fitting and feature embedding tailored to specific tasks. Secondly, KANs could be integrated into deep learning architectures by replacing the MLP components. Such integration may enhance KANs’ representation capacity in end-to-end optimization. Thirdly, the performance of KANs could be further assessed on other tasks, such as object recognition, to gain deeper insights into their strengths, limitations, and potential applications in feature representation.
@@ V Conclusions
In addition to SVR and MLP, six KANs are evaluated on four realistic databases for score prediction respectively using mid-level and high-level features. The results demonstrate that KANs achieve superior or competitive performance compared to SVR and MLP, highlighting their potential for enhancing performance in score prediction and related tasks.